1. The manager of the bank where you work tells you that your bank has $10.9...

1.

The manager of the bank where you work tells you that your bank has $10.9 million in excess reserves. She also tells you that the bank has $203 million in deposits and $164 million in loans. Given this information you find that the reserve requirement must be what percent?

Enter a number rounded to one decimal place. You may assume that the bank does not hold any other assets or liabilities besides reserves, loans, and deposits.

2.

Suppose we deposit $427 worth of currency into the bank. Banks make as many loans as possible, people hold no additional currency, and the reserve ratio is 10%. As a result of the loan process, what is the value of total deposits in the banking system? Enter a whole number with no dollar sign. Round to the nearest whole number.

3.

Bank 1 has deposits of $3,853 and reserves of $488. If the required reserve ratio is 10%, what is the value of the bank's excess reserves? Enter a whole number with no dollar sign. Round to the nearest whole number.

4.

Bank 1 has deposits of $4,456 and reserves of $706. If the required reserve ratio is 10%, what is the value of the bank's required reserves? Enter a whole number with no dollar sign. Round to the nearest whole number.
Answer #1

Answer1 ) We know that the Bank has deposits of $203m. The Bank has reserves of $10.9m

The Reserve Requirement is = 10.9/203 = 5.4%

The reserve requirement is the total amount of funds a bank must have on hand each night. It is a percentage of the bank's deposits. The nation's central bank sets the percentage rate.

Answer 2) We deposit 427 in the bank. The value of total deposits in the bank is 384 and 42 as reserve requirement (10%).

Answer 3) Bank has deposits of 3,853 and reserves of 488.

The bank is required to hold 10% of the deposits as reserve requirement. Hence it requires to hold 385.

however it is holding 488. So the excess reserve the bank is holding is 488- 385 = 103.

Answer 4) The bank holds deposits of 4,456 and reserves of 706. If the reserve ratio is 10%

the bank is suposed to hold 446 which is its required reserve.